please explain in detail 7

A particle with a charge of 16 μC experiences a force of 2.4×10−4 N when it moves at right angles to a magnetic field with a speed of 28 m/s . Part A What force does this particle experience when it moves with a speed of 6.4 m/s at an angle of 27 ∘ relative to the magnetic field? Express your answer using two significant figures. F = nothing N
Answer #1

as we know

F = qVB

2.4* 10^-4 = 16* 10^-6* 28 * B

B = 0.5357 T

now for an angle

F = 16* 10^-6* 6.4* 0.5357 sin 27

F = 2.5 * 10^-5 N

=======

Comment before rate in case any doubt, will reply for sure.. goodluck